HA and AH represents two 2-digit numbers. If HA - AH = 18, what is the value of the expressions H - A ?
Answer by htmentor(1343)

HA and AH represents two 2-digit numbers. If HA - AH = 18, what is the value of the expressions H - A ?
================================
The number HA can be written as 10H + A, since H represents the tens place and A the ones place
Similarly, AH = 10A + H
Since HA - AH = 18, we have
10H + A - (10A + H) = 18
9H - 9A = 18
9(H - A) = 18
H - A = 2